MONT 107N – Understanding Randomness Study/Discussion Questions on Primates and Philosophers February 23, 2010

Note: We will be discussing this text in class the week following Spring Break – March 8, 10, 12.

de Waal’s Essay “Morally Evolved”

1. Do you see humans as inherently social animals? Or are we inherently selfish loners who have banded together by a “rational” choice in order to maximize our individual advantage? The second point of view has been quite popular in some schools of thinking about politics, economics, etc. Have you run into this in courses you have taken in high school or here?

2. What is the “veneer theory” about human morality – the idea that de Waal attributes to T.H.Huxley? If morality is a “veneer,” what is underneath the veneer? How is this related to the Christian (especially Catholic) idea of original sin?

3. Are we good (when we are good) because of our nature and instincts? Or are we good because we fear punishment for “breaking the rules?”

4. What do Hume, Westermarck, etc. say about the role of the emotions in our moral choices? What are some examples they would use to make their point? Does this seem reasonable? If not, is our morality primarily a faculty connected with language, reason, and normative thinking (i.e. reasoning about what ought to be)? (Some famous philosophers, including for instance, Immanuel Kant, have certainly believed that opposing view!)

5. Does de Waal accept an important contribution from emotion in our sense of morality? What evidence does de Waal use to propose a basic continuity of the starting point for human ideas of morality with the social instincts and behaviors observed in higher primates? Does this convince you? And does this effectively demolish the “veneer theory” that de Waal is at pains to reject? Does Table 1 opposite page 23 seem a bit “stacked” in one direction?!

6. Why is it so easy for humans to sometimes behave terribly toward “others” – humans who are not members of their immediate circle of family, acquaintances, etc.? On the other hand, why are we at other times moved to acts of altruism (e.g. sending aid to Haiti to help earthquake victims to whom we may have no real connection at all)?

7. In his Appendix A, de Waal seems to be anticipating a possible criticism of his position. What would that criticism be and how serious do you think it is?

8. There was obviously some statistical analysis used to derive the plots like the one on page 47 in de Waal’s essay. What do you think the purpose of that was?

The Comments or Replies

1. What is the main point of Wright’s reply to de Waal? What does Wright think about the balance between emotion and reason in moral decisions?

2. Korsgaard is probably the respondent who is farthest from de Waal’s position – do you agree or disagree with this? In what ways does her point of view differ from his? What does she think are the distinctive features of human moral decision-making that make us different from animals. Why does she see humans as possibly “damaged?”

3. Kitcher’s criticism is quite “nuanced.” Can you summarize it? What is the “Hume-Smith lure” that he says de Waal and others have fallen into?

4. Singer discusses two different points. The first is a criticism of something he thinks de Waal “got wrong.” What is this? How serious a criticism do you think this is? The second responds to a point made in another of the appendices to de Waal’s essay. What did de Waal say about “animal rights?” What does Singer say? Does he agree or disagree with de Waal?

5. In his response to the responses, de Waal lays out in more detail what he thinks of as the aspects of morality, and how much of that is shared by primates and other non-human animals. What are his three levels of morality and how much of that hierarchy does he see in other animals?